13: A sample of carbon dioxide gas occupies 584 mL at 0.856 atm and 10 ◦C. What will the pressure be at a volume of 835 mL and a temperature of 24 ◦C? Answer in units of atm.
14: At 110◦C, the pressure of a sample of nitrogen is 1.78 atm. What will the pressure be at 267◦C, assuming constant volume? Answer in units of atm.
15: Suppose that 5.6 L of methane at a pressure of 876 Torr is transferred to a vessel of volume 4.1 L. What is the final pressure of methane if the change occurs at constant temperature? Answer in units of Torr.
1
Expert's answer
2020-04-19T07:09:13-0400

13). V1=584ml

P1=0.856atm

T1=283.15K

P2= ?

V2=835ml

T2=297.15K

"P1V1T2=P2V2T1"

P2=(0.856*584*297.15)/(835*283.15)

P2=0.628atm

14). T1=383.15K

P1=1.78atm

T2=540.15K

P2=?

"P1\/T1=P2\/T2"

P2=(1.78*540.15)/383.15

P2=2.51atm

15). V1=5.6L

P1=876 torr

V2=4.1L

P2=?

"P1V1=P2V2"

P2=(876*5.6)/4.1

P2= 1196.49 Torr